### v1

# Crafting Synthetic Realities: Examining Visual Realism and Misinformation Potential of Photorealistic AI-Generated Images

Abstract.
Advances in generative models have created Artificial Intelligence-Generated Images (AIGIs) nearly indistinguishable from real photographs. Leveraging a large corpus of 30,824 AIGIs collected from Instagram and Twitter, and combining quantitative content analysis with qualitative analysis, this study unpacks AI photorealism of AIGIs from four key dimensions, content, human, aesthetic, and production features. We find that photorealistic AIGIs often depict human figures, especially celebrities and politicians, with a high degree of surrealism and aesthetic professionalism, alongside a low degree of overt signals of AI production. This study is the first to empirically investigate photorealistic AIGIs across multiple platforms using a mixed-methods approach. Our findings provide important implications and insights for understanding visual misinformation and mitigating potential risks associated with photorealistic AIGIs. We also propose design recommendations to enhance the responsible use of AIGIs.
1. Introduction
The rise of AI-generated images (AIGIs) provides many opportunities for creative expression, but also raises concerns about the risks of misinformation through manipulated and synthetic visuals (Teachers College, Columbia University, 2023). In 2023, AI-generated images (AIGIs) of Donald Trump’s arrest and Pope Francis wearing a Balenciaga puffer jacket circulated widely on Twitter, deceiving many users (Devlin and Cheetham, 2023). On Reddit, an AI-generated image showing a blue plague in Russia, an event that has never happened, captured wide attention (Designboom, 2023). These incidents highlighted the risks and challenges of photorealistic AIGIs in misleading the public. News organizations increasingly face two major challenges brought by these AIGIs: avoiding deception by photorealistic AIGIs and accurately identifying and debunking misinformation (Royal United Services Institute, 2023).
AIGIs are produced entirely by artificial intelligence using algorithms that automate large-scale content production (Cao et al., 2023). Advances in generative models have created AIGIs that are nearly indistinguishable from real photographs with much lower cost and effort (Devlin and Cheetham, 2023). Hausken (Hausken, 2024) has introduced the concept of AI photorealism, which refers to the style of images that emulate photography using advanced AI tools. This study focuses on photorealistic AIGIs, and addresses the theoretical gap of conceptualizing AI photorealism along four feature-based dimensions of perceived realism: content features, human features, aesthetic features, and AI production features. Leveraging a large corpus of 30,824 AIGIs collected from Instagram and Twitter, we examined the visual features related to the perceived realism of these images. We developed a codebook with 42 categories to analyze AIGI features that fall into the four dimensions through content analysis. In addition, we conducted a qualitative analysis on human-related surrealism and aesthetic professionalism to understand high-level salient features.
Our research makes several contributions. First, our study comprehensively examines photorealistic AIGIs using a large corpus of real-world AIGIs collected from social media platforms. To our knowledge, this is the first study to empirically investigate photorealistic AIGIs (instead of artistic AIGIs) on multiple platforms (Twitter and Instagram) through a mixed-methods research design. Second, we provide a comprehensive theoretical framework to understand the AI photorealism of photorealistic AIGIs, which informs future research on the perception, discernment, and impact of photorealistic AIGIs. Third, we connect the features of photorealistic AIGIs uncovered in our analysis to their high potential for becoming misinformation through both direct and indirect ways. This provides essential insights for different stakeholders in AIGI production and consumption. We also propose design recommendations to better utilize and regulate AIGIs, which can inform better AIGI practices by social media platforms, AI companies, and creative industries.
2. Literature Review
2.1. AIGI and AI Photorealism
We begin our inquiry by distinguishing several intricately-connected concepts related to AI-generated visuals. First, photography is defined as the art or practice of using a camera to capture real scenes and moments (Hausken, 2024). With the principle that “seeing is believing” (Kasra et al., 2018) and the aim of providing evidential proof of events, photography captures reality directly through the camera lens, relying on the physical world, context, and timing.
However, photo-like images are not exclusively the product of photography. The progression of generative models has enabled AIGIs to mimic real photographs so closely that they often become indistinguishable from actual photographs, yet requiring less expertise and effort than ever (Devlin and Cheetham, 2023). Although these photorealistic AIGIs are not genuine photographs, they possess a high level of photorealism and blur the line between real and artificial in a more subtle way. This phenomenon is captured by the concept of AI photorealism (Hausken, 2024).
It is important to recognize that AI photorealism doesn’t emerge in isolation; it is rooted in the history of traditional photorealism predating generative AI. Photorealism involves creating images that simulate reality via artificial means. This concept has gained popularity and developed variants across various media. For example, in computer graphics, an image is considered photorealistic if it is generated by a computer yet indistinguishable from a photograph of the same scene (Ferwerda, 2003). In arts, photorealism is a genre that strives to create images so realistic they mimic actual photographs (Hausken, 2024). In mixed and augmented reality, photorealism is often pursued as an essential goal, characterized by visual coherence through the interaction of light, real objects, and virtual elements (Pereira et al., 2021). Despite these different definitions, one key element in defining photorealism is the deliberate attempt to replicate the appearance of reality, challenging perceptions by blurring the line between the real and the artificial. Research has shown that distinguishing between a photographic image and a photorealistic one is increasingly challenging, even for experts (Lehmuskallio et al., 2019; Shen et al., 2021), as audiences are likely to believe that images are authentic due to their truth default (Levine, 2014).
The emergence of AI tools such as MidJourney and DALL-E enables the creation of photorealistic visuals at scale. This complicates the definition of photorealism as these AI technologies far surpass previous methods in creating visuals with greater detail and much higher levels of realism. Hausken (Hausken, 2024) introduced the concept of AI photorealism, a style of images that emulate photography using advanced AI tools, which differs from both traditional photography and the conventional concept of photorealism, as shown in Table 1. However, a comprehensive and systematic conceptualization of AI photorealism is still lacking.
In this paper, we focus on photorealistic AIGIs and further unpack the concept of AI photorealism. While photorealistic AIGIs hold broad implications for creative industries, their proliferation, particularly in the form of highly realistic but false images, can accelerate the spread of misinformation and undermine trust in news media. Thus, this paper looks into the specific features of AI photorealism within photorealistic AIGIs and addresses the implications of these features in the context of misinformation.
| Definition | |
| Photography | The art or practice of using a camera to capture real scenes and moments, relying on the physical world, context, and timing. |
| Traditional Photorealism | A genre of art that strives to create images so realistic they mimic actual photographs. |
| AI Photorealism | The development of photorealism in the age of AI — the style of images that emulate photography using advanced AI tools, with images created entirely by AI algorithms. |
Definitions of Photography, Traditional Photorealism, and AI Photorealism
2.2. Unpacking AI Photorealism through the Lens of Perceived Realism
To understand what features contribute to higher AI photorealism in AIGIs, we draw from the concept of perceived realism, defined as the extent to which an image corresponds to the original scene from the viewer’s perspective (De Ridder, 1996; Lee et al., 2013). Previous studies in computer vision, graphics, and visualization have explored the visual features — their presentation, existence, and combination — that lead people to consider as visual realistic. These studies have explored various forms of visuals, including real-life photos, computer-generated images, mixed reality simulations, and most recently, AIGIs (Chen et al., 2024; Fan et al., 2014, 2020; Lee et al., 2013; Lu et al., 2023; Rademacher et al., 2001; Wang et al., 2024).
In this study, we categorize the visual features related to the perceived realism of AIGIs into four dimensions: content features, human features, aesthetic features, and AI production features. We explain each dimension in detail below.
2.2.1. Content Features
Content features refer to the visual presentation of objects within the image, including their presence and combinations, such as text captions, humans, and animals (Peng et al., 2023a; Qian et al., 2024). These features demonstrate how contextually accurate and believable the depiction of objects appears. Empirical evidence suggests that the number of unique objects, their natural appearance, and their combinations significantly influence perceptions of visual realism (Fan et al., 2014, 2020). For example, Fan et al. (2014) found that the natural combination of objects positively correlates with perceived realism, while the number of objects negatively correlates with it. Similarly, Rademacher et al. (2001) found that perceived realism decreased when the number of objects increased with a borderline-significant correlation. Building upon these studies, we focus on different types of objects and their combinations as our content features.
2.2.2. Human Features
Humans, as a specific category of objects, have significant implications for visual engagement and perceived realism. Previous research finds that the presence of humans and human faces may increase the engagement of social media posts (Li and Xie, 2020; Lu and Peng, 2024; Peng, 2021). Facial features, such as presentation of faces and identities also influence social media engagement (Bakhshi et al., 2014; Joo et al., 2014; Muñoz and Towner, 2017). Although research on the relationship between human faces and perceived realism of images is still limited, one study on computer-generated faces found that the presence of complete faces triggers higher realism perception than partial faces (Fan et al., 2014). In this study, we focus on the presence of human faces, their identities, and the interaction between human and other elements to derive more comprehensive features of humans.
2.2.3. Aesthetic Features
Visual aesthetics, which influences human perceptions of visual attributes (Peng, 2022; Pieters, 2010) and diffusion of visuals (Bakhshi and Gilbert, 2015; Sharma and Peng, 2024), have been extensively explored in prior research using various indicators, such as colors, composition, visual complexity, and quality (Kanuri et al., 2024; Lu and, Cindy; Peng, 2022; Peng et al., 2023a; Qian et al., 2024). In the assessment of AIGIs and other visuals, color and quality are two crucial factors in considering visual realism (Chen et al., 2024; Cutzu et al., 2003; Lu et al., 2023). Wang and Doube (2011) built realism metrics using color variance, shadow softness, and surface roughness to evaluate visual realism in video games. We build on these studies to capture both low-level features such as color and image quality, as well as higher-level aesthetic combinations as aesthetic features.
2.2.4. AI Production Features
AI production features reflect the development quality of the generation tool and can predict the audience’s ability to identify AIGIs. Flaws, such as irrational or illogical contents, contradictory elements, or content discontinuity, are often considered AI production features and are categorized under “rationality” or “artifacts” in previous studies of visual realism (Chen et al., 2023; Lu et al., 2023). Additionally, AI production quality can influence perceptions of realism. Unlike aesthetic quality, which pertains to resolution and composition, AI can sometimes produce high-resolution images with low-quality details, such as showing six fingers on a person’s hand in an otherwise high-aesthetic image. Compared with the traditional understanding of perceived realism, these production features are uniquely specific to the AI-generation context.
Grounded in theories related to these four dimensions, the following research questions were proposed to understand features that contribute to AIGIs’ perceived realism:
RQ1: What contents are present in photorealistic AIGIs?
RQ2: What human features are present in photorealistic AIGIs?
RQ3: What aesthetic attributes are present in photorealistic AIGIs?
RQ4: What AI production features are present in photorealistic AIGIs?
Flow Chart of Data Collection and Analytical Methods. The research design chart illustrates the process of data collection, codebook development, and analysis for a study on viral photorealistic AI-generated images. Data is collected from Instagram and Twitter accounts using snowball sampling, yielding over 30,000 images. These images are sampled and categorized into a codebook through both quantitative and qualitative analyses. Quantitative analysis involves coder training and independent coding, while qualitative analysis focuses on recording and discussing salient features of the images.
| Variables | Definition | Cohen’s Kappa |
| Content features | ||
| 1. Human | The presence of an object that is human being | .91 |
| 2. Human-like | The presence of an object that has the shape of human-beings but not human or hard to be categorized as human, e.g., ghosts, aliens, human-shape robots etc. | .68 |
| 3. Non-human living beings | The presence of an object that has life but is not a human or human-like object. | .84 |
| 4. Non-living objects | The presence of an object that has no life. | .96* |
| 5. Add-on captions | The presence of a caption with understandable content that deliver a specific message, yet are an addition to the image and do not inherently belong to the image itself. | .97* |
| 6. Fictional captions | The presence of a caption composed of fictional characters or letters, representing languages that do not exist in the real world, thus making their meanings unidentifiable. | .83 |
| 7. Other captions | The presence of a non-distorted in-image caption. | .68 |
| 8. Disaster | The presence of an event or situation that causes significant and widespread damage, destruction, disruption, or distress, often resulting in adverse effects on human life, property, the environment, and the overall well-being of a community or society; scalable, which means it need to be applied to a large number of people or to buildings. | 1.00* |
| 9. Surrealism combination | The presence of unnatural pairings of situations, human, or objects not typically observed together in reality. | .79 |
| Human features | ||
| 1. Face | More than 50% of the face is identifiable and is clear and large enough for viewers to detect its identity or emotions. | .96 |
| 1.1 Identity | Human identity is identifiable based on the presented human face. | 1.00* |
| 1.1.1 Politician | The presence of any politicians that can be recognized based on the best of knowledge. Politicians refer to people who work in government, making rules and decisions that affect the country or region they represent. | 1.00* |
| 1.1.2 Celebrity | The presence of any celebrities who are not political figures and can be recognized based on the best of knowledge. | 1.00* |
| Aesthetic features | ||
| 1. Colored images | If it is a colored image. | .85 |
| 2. High image quality | The image is clear, with high resolution. | .97* |
| Production features | ||
| 1. Production flaw | The presence of any obvious production flaws can be told from unequivocal evidence that this image is made by AI. | .76 |
| 2. Watermark | The presence of a watermark that explicitly indicate the image is AI-generated. | .66 |
| Note: * indicates that we used percentage agreement instead of Cohen’s Kappa because the skewed distribution of our 0/1 coding can make Cohen’s Kappa less reliable. When one category is infrequent, kappa values can be misleadingly high or low. | ||
Variables, Definitions, and Cohen’s Kappa of Each Variable.
3. Data and Methods
3.1. Social Media Data
We investigate AIGI features through a comprehensive analysis of 30,824 AIGIs collected from Twitter and Instagram, as both platforms have substantial user bases where AIGIs are widely shared and engaged with. According to Pew Research Center (Pew Research Center, 2023), 62% of U.S. teens use Instagram to share images, while 69% of Twitter users regularly using the platform for news and real-time updates. We have also seen viral AIGIs used for misinformation on these platforms (Chen and Zou, 2024; Devlin and Cheetham, 2023; PetaPixel, 2023).
We gathered data through a two-step process, as Figure 1 shows. First, we identified AIGI accounts that frequently posted AI-generated images from July 2022 to August 2023. We started from examining news articles mentioning viral photorealistic AIGIs to identify seed accounts on Twitter or Instagram where these AIGIs originated from. Then, we snowballed and collected more AIGI accounts that were mentioned, forwarded, or algorithmically recommended by those seed accounts. We included AIGI accounts where at least 50% of the images posted were photorealistic AI-generated images. We terminated the sampling process once we consistently observed repeated mentions, forwards, and recommendations involving the same accounts. Ultimately, we included 49 Instagram accounts and 60 Twitter accounts.
From these accounts, we collected images from their timeline posts between July 12, 2022, and August 31, 2023. The start date was chosen to align with the public release of the most widely used AI image-generating tools, Midjourney and DALL-E. The end date was set for consistency, as our image collection began in September 2023, and we aimed to standardize the end time across all accounts. For Instagram, we utilized the 4k Stogram1114K Stogram is a desktop application for downloading photos, videos, and stories from Instagram using usernames, hashtags, or locations. https://www.4kdownload.com/products/stogram-8 tool to download the timeline posts and their images, accumulating a total of 28,412 images, from which we randomly selected 2,000 images for analysis. To match the Instagram sample, we manually downloaded approximately 60 images from each Twitter account during the same time period as we did for Instagram, resulting in a total of 2,412 images.222Due to a change in the Twitter/X API, we were unable to easily automate the retrieval of a large number of images from this platform. As a result, we resorted to manually downloading them. In addition, during the image scraping process, two Instagram accounts and two Twitter accounts were excluded due to access issues. We did not apply any manual filtering to ensure consistency between the platforms, the only exclusions being screenshots that were clearly not AI-generated.
3.2. Analytical Methods
We started our analysis from a quantitative content analysis. We developed a codebook to analyze AIGI features related to perceived realism using an exploratory and iterative approach. We designed the codebook firstly by drawing from popular examples of AI-generated images that went viral on social media, mislead people, and attracted sizable media attention. This approach allowed us to focus on the types of AIGIs most likely to mislead audiences. Then, one author read a random sample of AIGIs during the data collection phase, recorded the features, and drafted the initial version of the codebook. Two additional authors independently reviewed the collected images based on the draft and provided feedback. Then, all six authors collaboratively refined and finalized the codebook. This process was repeated until no significant refinements were made to the codebook.
The codebook starts by identifying whether the AIGI is photorealistic, which is crucial for filtering out AIGIs that were not reflecting reality. After excluding non-photorealistic images, 1,785 images from Instagram and 1,858 images from Twitter were included for further analysis.
The codebook includes 42 categories, 15 of which are first-order variables coded for each image, while the remaining 27 are second-order variables nested within selected first-order categories. All variables are binary variables. Corresponding to the four theoretical dimensions of perceived realism in AIGIs, the major variables can be seen in Table 2. The first section “Content Features” focuses on the elements depicted in the images, such as textual captions, presence of humans, non-human living beings, and non-living objects. In addition, as AI-generated disaster images have been circulating on social media (Designboom, 2023), potentially misleading the public and evoking fear or sympathy (Wang et al., 2021; Gupta et al., 2013), we included the portrayal of disasters and plagues. We also captured the surreal combinations — unnatural pairings of situations, human, or objects not typically observed together in reality, are defined by us with roots in the art movement, where surrealism is characterized by juxtaposing unrelated, dream-like elements, people, or situations (Germain, 1974). The second section, “Human Features,” includes the presence of human faces and distinguishes identifiable individuals, particularly politicians and celebrities. The third section, “Aesthetic Features,” encompasses aspects like color schemes and image quality. Finally, the fourth section captures “AI Production Features” such as production flaws and watermarks.
Two trained coders independently coded a random sample of 200 AIGIs. Discrepancies were discussed until a satisfactory level of agreement was achieved. All variables demonstrated acceptable intercoder reliability. Following content analysis, we conducted a descriptive analysis using R. We separated images from two platforms for analysis and conducted a series of chi-square tests and independent t-tests to compare the results, as shown in Appendix A.1
In addition to quantitative analysis, we conducted qualitative analysis on two high-level salient features emerging from both theory and our collected data—human-related surrealism and aesthetic professionalism. To better understand and categorize these features, we assessed a sample of images exhibiting each salient feature, using an open coding strategy. After discussing these observations among all authors, we derived more granular subcategories within each feature.
Percentages of AIGI Visual Features. The bar chart displays the distribution of various features in AIGIs on Instagram and Twitter, categorized by content, human-related features, aesthetic qualities, and production attributes. The chart includes comparisons across both platforms, with percentages representing the presence of each feature. Key features include content types such as non-living objects, humans, and surrealism combinations, as well as human-related features like faces and celebrity identities. Aesthetic features cover colored images and image quality, while production flaws and watermarks represent production characteristics. For some features, including face and identity-related features, their percentage is calculated based on the total number of their antecedent features.
4. Results
4.1. Quantitative Analysis of AIGI Features
Figure 2 shows the descriptive results of AIGI features in four perceived-realism dimensions, which indicate that AIGIs exhibit distinct visual features.
4.1.1. Content Features
Photorealistic AIGIs depict a wide variety of objects. 98.4% of photorealistic AIGIs on Twitter and 98.6% on Instagram contain non-living objects such as everyday buildings, vehicles, and landmarks such as Eiffel tower. AIGIs present a strong emphasis on human subjects, with figures of 67.5% on Twitter and 68.7% on Instagram. In addition, 28.2% on Twitter and 26.6% on Instagram feature non-human living beings such as real-life animals and fantastical creatures like monsters, which can blur the line between reality and fabrication. The mix of real and fictional elements can make it difficult for viewers to discern factual content from misleading or deceptive images, as people may misinterpret these visuals as representing real events or entities. Surrealism combination was also prevalent in AIGIs, at 58.2% on Twitter and 60.8% on Instagram, reflecting substantial incorporation of imaginative and unreal combinations in the visuals. Chi-square tests confirmed similar patterns for content features across the platforms, with minor variations, as detailed in Appendix A.1.
4.1.2. Human Features
Through a further analysis on images containing humans, we find that 73.4% of Instagram images and 71.0% of Twitter images displayed human faces. Among these, more than 10% were identifiable as politicians or celebrities, with significant platform differences. Instagram featured more politicians, such as Donald Trump, Barack Obama, and Queen Elizabeth II, while Twitter displayed more celebrities, such as Elon Musk, Mark Zuckerberg, and Kanye West.
4.1.3. Aesthetic Features
Photorealistic AIGIs exhibit high levels of aesthetic features on both Instagram and Twitter, with 98.0% of Twitter photorealistic AIGIs and 93.0% of Instagram photorealistic AIGIs being colored images. Additionally, over 92% of photorealistic AIGIs on both platforms display high image quality.
4.1.4. AI Production Features
Compared to the prevalence of content, human, and aesthetic variables, photorealistic AIGIs on Twitter and Instagram exhibit a low level of AI production features. Approximately 8% of photorealistic AIGIs on both platforms contain observable production flaws, which appear in various forms. For example, some AIGIs present strange facial or body features, such as misaligned eyes, disproportionate mouths, and twisted limbs where arms or legs bend at impossible angles. There are also instances of incorrect finger placement, where a hand may have an unnatural number or arrangement of fingers. Blurriness is another common issue, particularly in areas that should be sharp, such as the outlines of objects or facial features, giving the image an artificially processed look. Additionally, some images exhibit mismatched borders, where the edges of objects or people don’t align properly, indicating poor rendering or stitching in the AI generation process. Despite these production flaws, only 1.5% of Instagram AIGIs contain watermarks signaling AI production, while this figure is lower on Twitter, at 0.8%.
4.2. Human-related Surrealism
We then conducted a qualitative analysis to gain a deeper understanding of the AIGIs in our sample, focusing on two salient features: human-related surrealism and aesthetic professionalism. Through our qualitative analyses of human-related surrealism in photorealistic AIGIs, we find that AIGIs depict humans in surrealistic settings by manipulating one or more of the following elements: 1) physical characteristics, 2) behaviors and interactions, and 3) contexts.
4.2.1. Surreal Physical Characteristics
We find a large body of human-based surreal AIGIs portraying a single person, which manipulate the person’s physical characteristics, including facial abnormalities, body abnormalities, and clothing incongruence. For facial abnormalities, AIGIs usually present a distorted or asymmetric face, as demonstrated in Figure 3 Panel 1A, or an exaggerated expression with unlikely intensity. For body abnormalities, AIGIs usually present incomplete bodies or over-engineered functional body features, as Panel 2A shows, a woman’s upper body is shown in a bakery. Besides physical facial and body abnormalities, AIGIs may also present clothing or costumes that are normally not seen in real life, including unusual clothing segments and accessories, or not commonly expected, as Panel 3A shows.
AIGI examples of human-based surrealism. The image presents AI-generated visuals divided into four categories: A depicts surreal physical characteristics of ordinary individuals, B features surreal physical characteristics of politicians and celebrities, C shows surreal behaviors and interactions, and D portrays surreal contexts. Each category contains three representative images with exaggerated or surreal elements, highlighting the creative manipulation of appearances, actions, and environments.
Among all the human beings depicted in AIGIs, people with recognizable identities play a particularly important role in presenting surrealistic combinations. Politicians are highly likely to have their physical characteristics manipulated in surreal AIGIs. Our dataset reveals a salient presence of the former U.S. Presidents such as Donald J. Trump and Barack Obama, whose physical features were surrealistically altered. One such image, shown in Figure 3 Panel 1B, portrays Barack Obama in a leather jacket at a public protest or concert, where his facial expression of excitement is highly exaggerated. This resonates with prior research suggesting that images of politicians in non-political settings and displaying emotions may elicit more online engagement (Peng, 2021). Another portrayal of Obama in Figure 2B, features him wearing indigenous attire, including necklaces and a chest ornament, which deviates from the traditional attire typically worn by Obama.
In addition to politicians, celebrities from various fields, including actors and businessmen, also frequently appear in AIGIs featuring surreal physical characteristics. Elon Musk, a globally renowned entrepreneur, is prominently featured in these AIGIs. As shown in Panel 3B, the excessively large muscles and the absence of natural body textures exemplify the surrealism in Musk’s visual representation. Moreover, combining his formidable physique with his calm and serious facial expression, the image portrays a heightened sense of masculinity and a strong-man image of Musk.
4.2.2. Surreal Behaviors and Interactions
Human-related surrealism is also reflected in the portrayal of human behaviors and interactions. This type of surrealism portrays either supernatural behaviors or normal behaviors in unrealistic or even bizarre occasions. For example, Panel 1C in Figure 3 shows a group of people clumping together and floating in mid-air, which violates the law of physics. The calm and normal reactions shown on the faces of both the floating people and the bystanders further enhance the surrealism of this image.
Politicians and celebrities are commonly depicted with surreal behaviors and interactions in AIGIs. In images that portray a single figure, surreal behaviors emerge through the juxtaposition of a powerful figure and an unexpected action. For example, Panel 2C shows Pope Francis engaging in skateboarding, a highly unlikely activity given his physical condition and role as a religious leader. Surrealism also occurs when depicting multiple politicians or celebrities in their interactions. Panel 3C depicts Donald Trump and Queen Elizabeth II sharing a kiss, which is both implausible and absurd. We found similar intimate behaviors depicted in AIGIs that involve other politicians and celebrities, such as Donald Trump and Joe Biden, as well as Elon Musk and Mark Zuckerberg.
4.2.3. Surreal Contexts
Finally, surrealism occurs when humans are placed in nearly impossible or supernatural contexts. As Figure 3 Panel 1D shows, an elderly woman in a traditional outfit with a joyful expression is accompanied by a savage-looking dragon. The presence of this fantastical beast conveys surrealism by combining ordinary and extraordinary elements.
In addition to ordinary people, we also find AIGIs that place politicians and celebrities in surreal contexts. For example, Panel 2D shows Elon Musk, in a formal but burnt suit, standing confidently in a close distance with flames and destruction, which is also impossible in real life. The mixture of elements that represent order and elegance (suit, rose, calm expression) with those that represent chaos (flame, destruction, burnt clothes) exemplifies surrealism and may portray a masculine and authoritative image of Musk. Groups of individuals can also serve as surreal contexts. In Panel 3D, Donald Trump is surrounded by people dressed in Star Trek-style uniforms, with incongruent head sizes, positions, and some in bizarre outfits. This depiction of surrounding individuals seems more surreal in the spaceship-like background.
4.3. Aesthetic Professionalism
Our collected AIGIs also reflect a high level of aesthetic professionalism, constructed by the staging and professional-quality lightening. Staging, or staged photography, manipulates the subject with controlled theatrical effects, lighting, and settings in photography (Castro Díaz, 2019; Neumüller, 2018). As Panel A in Figure 4 shows, the model was arranged in a deliberate position, with a fashionable or exaggerated costume, a pure-color background, and crafted studio lighting. Products and nonhuman objects can also be the subject of a staged photograph. Panel B represents a staged photograph of a sports car, positioned in an outdoor setting with artificial lighting and at a particular angle. The car’s shiny exterior, crafted paint, and sleek finishes make it more luxurious and visually appealing.
Staging and professionalism examples. The image displays eight AI-generated visuals labeled A through H, demonstrating a range of staging and professionalism. These include a fashionable figure in a wide-brimmed hat (A), a futuristic sports car (B), a stylized portrait of a person in a hoodie (C), a cyberpunk character (D), a woman in an industrial setting (E), an older woman smiling in a red headscarf (F), an older woman indoors in a contemplative pose (G), and a man in a suit holding a gun, resembling a spy or action figure (H).
The staging feature corresponds to how AIGIs generally follow conventional photographic standards of professional quality. For example, Panels C and D in Figure 4 tend to have high resolution and vivid detail. Notably, many of the AIGIs have one focal subject and a shallow depth of field, as shown in Panels E and F. This refers to a photographic technique where only a small portion of the image is in sharp focus, while the rest appears blurred. The primary subject, usually in the foreground, remains sharp and clear, while the background and sometimes even the foreground elements become out of focus. This technique is often used to draw attention to the main subject and minimize distractions. In photography, this effect is typically achieved by using a large aperture on the camera lens. Recently, some smartphone cameras have been designed with algorithms to mimic this effect as well. Nonetheless, a large percentage of AIGIs follow this convention, demonstrating signs of professional quality.
Another indicator of aesthetic professionalism in photorealistic AIGIs is the expert use of lighting. For instance, in images like Figure 4 Panels G and H, the light is well-arranged, often lit from one side to create a dramatic effect of light and shadow in portraits. One image may show the use of “Rembrandt lighting,” a technique that creates a small, triangular patch of light on the cheek of the subject, while half of the face is lit and the other half is in shadow. These dramatic lighting effects are characteristic of professional portrait photography, suggesting that these AIGIs adhere to high standards of visual aesthetics.
5. Discussion
This study, one of the first thorough examinations of photorealistic AIGIs circulated on different social media platforms, empirically investigates the visual features that contribute to perceived realism and the potential implications of these realistic depictions. Grounded in the theory of photorealism and perceived realism, we conceptualize AIGI photorealism into four dimensions: content features, human features, aesthetic features, and AI production features. Through a comprehensive analysis of 30,824 AIGIs from Twitter and Instagram using quantitative content analysis and qualitative image analysis, we identify the following four key patterns corresponding to these dimensions.
First, among a diverse range of object depictions, there is prevalent presence of human figures among photorealistic AIGIs, with over 67% containing humans and approximately 10% of these human-related images featuring politicians and celebrities. Second, there is a notable presence of human-related surrealism, where AI manipulates physical characteristics, behaviors, interactions, or contexts to create visually striking and often distorted images. Additionally, the AIGIs consistently exhibit a ubiquitous color scheme, high image resolutions, and aesthetic professionalism conveyed through staging and professional lighting. Finally, despite the advanced nature of AIGIs depicting realistic scenarios, only a small proportion show overt signs of AI generation, such as production flaws or watermarks, making them nearly indistinguishable from human-created visuals.
5.1. Theoretical Implications
This study extends the theory of photorealism in the age of artificial intelligence through the lens of perceived realism with a four-dimensional framework. Building upon prior work on photorealism (Ferwerda, 2003; Hausken, 2024), this study addresses a theoretical gap by conceptualizing AI photorealism through an empirical approach. The framework includes theoretically-grounded features that are predictive of human perceptions and engagement with visuals, such as content, human, and aesthetic dimensions, as well as an AI production feature that is particularly relevant in the AI context. The imbalanced presentation of these features highlights the need for future research to empirically test their impact.
Our findings that AIGIs often mix realistic imagery with surreal content, particularly in depictions of human figures such as politicians and celebrities, also challenge existing theories of perceived realism and authenticity. According to the Truth Default Theory (TDT), human cognition tends to default to trust, minimizing the cognitive cost of questioning information authenticity (Levine, 2014). AIGIs may exploit this trust by seamlessly merging real and fictional elements in high-quality visuals, leading viewers to overlook surrealistic or impossible aspects. This issue becomes particularly problematic when surrealism is integrated into realistic contexts without overt contradictions. The surreal imagery can capture more attention by presenting unusual or illogical elements that deviate from expectations (Itti and Koch, 2000), potentially leading to memory confusion, where viewers might later recall these fabricated images as real. At the same time, the use of sophisticated aesthetic appeals and techniques in manipulated photography also highlights the manipulative power of AIGIs, especially in combination with surrealism portrayals. Thus, it is crucial to consider the potential harm and practical implications regulating the use of AIGIs, which we focus on their potential use in misinformation and potential regulations.
5.2. High Misinformation Potentials of AIGIs
Collectively, the features of photorealistic AIGIs uncovered in our analysis point to their very high potential to become misinformation — either intentionally through disinformation campaigns or inadvertently as they spread organically via online platforms, out of their intended context. First, the prevalence of humans, especially politicians and celebrities, coupled with a highly professional aesthetic style, provides the perfect recipe for mimicking photorealistic images usually found in news media. For instance, the surrealism presentation of political figures may negatively influence public’s perceptions and trust in political figures by “kicking up” (Wagg, 1998). Second, the AIGIs we analyzed all lack discernible AI watermarks, therefore making it difficult for online users to trace where these images are from, especially when such images would be placed in a misleading context or blend with a high level of surrealism. Third, previous research has shown consistently that the presence of human faces (Li and Xie, 2020; Lu and Peng, 2024) and high production quality (Li and Xie, 2020), features we found in AIGIs, contribute to higher user engagement such as likes and retweets. In other words, the photorealistic AIGIs we analyzed not only look real and resemble the images in credible news reports in both content and style, but they are also likely to engage online users and potentially reach a large audience.
In addition to becoming misinformation directly, photorealistic AIGIs may indirectly contribute to the spread of misinformation by providing clickbait or sensational content to online misinformation sites, sometimes known as “content farms” or “click farms” (DiResta and Goldstein, 2024). This type of operation is essentially based on a bait-and-switch tactic that leverages highly sensational or entertaining content to grow their audience base and boost engagement, before disseminating misleading information to the audience network (Peng et al., 2023b). Generative AI has dramatically reduced the cost to produce the highly viral content needed to build a large following, and the surrealism element we discovered in our analysis could significantly contribute to the visuals’ novelty, virality and entertainment value. For example, a well-documented case of such AI-generated visuals is “Shrimp Jesus” (DiResta and Goldstein, 2024), which is emblematic of a known figure (Jesus) with surreal physical characteristics and contexts, two of the surrealism formats we discovered here. While such surreal AIGIs may not be perceived by online users as a direct manifestation of reality, they nonetheless draw users’ attention and engagement, leading them to scams and disinformation campaigns. Taken together, we argue that the direct and indirect misinformation potential of photorealistic AIGIs cannot be overestimated, and it is imperative to build resilience against such potential.
5.3. Toward Better Use and Regulation of Photorealistic AIGIs
We emphasize the following design recommendations to improve the use and regulation of AIGIs, especially photorealistic AIGIs. First, in our study, only a small percentage of images displayed noticeable AI production flaws, and these often appeared in subtle ways that human coders would miss unless they were paying very close attention. Moreover, the flaws we observed were diverse and varied, requiring viewers to examine various aspects of the images. These findings may imply that AI-literacy interventions aimed at addressing production flaws may not be ideal. For example, AI literacy programs designed to teach average social media users how to identify such flaws or obvious AIGI indicators (Malcolm, 2024) may not be as effective as expected. Similarly, algorithms developed to detect these flaws as markers of AI-generated images also face limitations. Interventions centered on detecting AI production flaws may soon become outdated as generative AI technologies evolve rapidly, while also requiring intensive human attention that is often impractical. While these approaches might offer short-term solutions for individual users, they are not viable in the long run.
Therefore, we argue that addressing AI-generated visual misinformation should not be the sole responsibility of individual users, but rather a shared responsibility at a broader level, involving AI developers, social media companies, and other key stakeholders. For example, social media platforms and AI companies must prioritize transparency by incorporating features that help users easily distinguish photorealistic AIGIs from real photos. Our study finds that the vast majority of AIGIs are not labeled. Today, some platforms, like Instagram, allow users to include AI-generated labels when uploading their images (Instagram, 2024). However, these disclaimers or AI labels can easily disappear when images are reshared on other platforms given that these AI indicators are not embedded in the images themselves. A potential solution is the implementation of mandatory labels or other provenance cues directly embedded in the content. These cues can be integrated within the algorithm itself, which ensures that the labels remain intact across different platforms. Such markers would provide an immediate visual cue to users, indicating that the content they are viewing was generated by AI, thereby increasing discernment and reducing the risk of misinformation. By clarifying the origins of digital content, these platforms can cultivate a more informed and critically engaged user base.
In addition, our study highlights the critical need to closely monitor and regulate AI-generated content (AIGCs) that involves human subjects, particularly when it includes identifiable faces of public figures such as politicians and celebrities. Our research found a significant percentage of the images featuring human faces, many of which were of well-known individuals. These images are especially vulnerable to misuse, as they can easily be manipulated to spread misinformation or create deceptive narratives toward the public figures. The recent misuses of ordinary people’s faces in deepfake pornography and criminal activities (Mackenzie, 2024) elevate the need for regulation in using ordinary people’s faces in AIGIs. AI and social media companies should implement strict rules that govern the inclusion of real person’s images in AI-generated content, ensuring that such usage does not contribute to the dissemination of false information or harm the reputations of those depicted.
Finally, our study provides insights into the implications of AI-generated content for creative industries. One notable observation is that the AIGCs in our sample are mimicking the high-quality features of professional photography, such as staging, expert lighting, and composition. Traditionally, the expertise and unique vision of human artists have been crucial to make high-quality images. Yet, the ability of AI to replicate these professional qualities suggests that AIGCs could be increasingly used in commercial contexts, potentially reducing the demand for human photographers and artists. On the other hand, the rapid development of AIGCs also opens up new possibilities for design and creativity. Designers and content creators might leverage AI tools to experiment with new ideas, enhance their work, or produce content at a scale that would be impossible to achieve manually. Despite these opportunities and changes, practices that protect copyright in creative industries remain necessary.
5.4. Limitations
This study has several limitations. First, our analysis was limited to a select number of accounts on Instagram and Twitter, which may not provide a comprehensive understanding of the overall patterns and trends in photorealistic AIGIs across all social media platforms. This limited scope may overlook less active accounts or other platforms where AIGIs are prevalent, such as Facebook or TikTok.
Additionally, this study focused exclusively on visual content without analyzing the accompanying textual elements, such as captions or embedded text within the images. These textual elements play a significant role in shaping the interpretation and engagement with visual content, and their exclusion may result in a less comprehensive analysis.
Moreover, this study concentrated on photorealistic AIGIs without considering other genres or styles of AI-generated content that might influence public perception differently. This focus might limit the generalizability of our findings to other types of AI-generated visuals. Furthermore, the descriptive nature of this study cannot afford capturing how these images impact viewer trust and perception, especially in diverse cultural contexts.
5.5. Future Directions
To expand the current research focus, we propose several promising directions for future research. First, integrating textual analysis of captions and embedded text with visual content analysis, as well as considering other contextual social media features, will provide a more holistic understanding of how AIGIs are presented and interpreted. Prior research has shown that nonvisual features, such as the original source and virality metrics (number of likes and shares), significantly influence credibility perceptions (Sundar and Shyam, 2008). User characteristics, such as demographics, digital media literacy, and prior experience with digital imaging, also affect their susceptibility to misinformation (Shen et al., 2019). Future studies can explore how visual features in photorealistic AIGIs interact with these nonvisual elements and how they function differently across various user groups.
Second, future studies can include a broader range of social media platforms and a larger number of accounts to capture a more comprehensive view of AIGI trends and patterns. Expanding the scope to include various platforms, accounts, and genres of AI-generated content across different platforms can help determine whether the identified features and their impacts on perception are consistent across different types of AI-generated visuals.
In addition, this research can inform future research in the realm of visual misinformation (Peng et al., 2023a). To pinpoint the visual attributes that influence human judgment of realism and credibility of visual misinformation in an AIGI context, machine-learning-based feature selection methods, such as LASSO, can be used to identify important factors (Matz et al., 2019). Researchers can also use models to understand the relationships among attributes (Iigaya et al., 2021), clustering algorithms to identify intercorrelated features (Huang, 1997; Lu and, Cindy), or activation maps to highlight important image regions that influence perceptions (Joo and Steinert-Threlkeld, 2022). By pursuing these research directions, future studies can provide a more comprehensive understanding of AIGIs and their impact on viewer perceptions, ultimately contributing to more effective management and mitigation of misinformation in various media environments.
Acknowledgements.
This work was partially supported by the National Science Foundation (CNS-2150716, CNS-2150723). We thank Ziqi Liu for superb research assistance. We also thank members of the Computational Multi-Modal Communication (CMMC) Lab and participants of the 74th Annual Conference of the International Communication Association Images-as-Data panel for their valuable feedback.References
- (1)
- Bakhshi and Gilbert (2015) Saeideh Bakhshi and Eric Gilbert. 2015. Red, Purple and Pink: The Colors of Diffusion on Pinterest. PLOS ONE 10, 2 (Feb. 2015), e0117148. https://doi.org/10.1371/journal.pone.0117148
- Bakhshi et al. (2014) Saeideh Bakhshi, David A. Shamma, and Eric Gilbert. 2014. Faces engage us: photos with faces attract more likes and comments on Instagram. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. ACM, Toronto Ontario Canada, 965–974. https://doi.org/10.1145/2556288.2557403
- Cao et al. (2023) Yihan Cao, Siyu Li, Yixin Liu, Zhiling Yan, Yutong Dai, Philip S. Yu, and Lichao Sun. 2023. A Comprehensive Survey of AI-Generated Content (AIGC): A History of Generative AI from GAN to ChatGPT. https://doi.org/10.48550/ARXIV.2303.04226 Version Number: 1.
- Castro Díaz (2019) Maribel Castro Díaz. 2019. Visual Storytelling in Hypermodernity: The Transformative Construction of Symbolic Realities Through Staged Photography. Journal of Comparative Literature and Aesthetics 42 (Nov. 2019), 84–99.
- Chen et al. (2024) Jian Chen, Petra Isenberg, Robert S. Laramee, Tobias Isenberg, Michael Sedlmair, Torsten Moeller, and Rui Li. 2024. An Image-based Typology for Visualization. https://doi.org/10.48550/ARXIV.2403.05594 Version Number: 2.
- Chen and Zou (2024) Yiqun Chen and James Y Zou. 2024. Twigma: A dataset of ai-generated images with metadata from twitter. Advances in Neural Information Processing Systems 36 (2024), 37748–37760.
- Chen et al. (2023) Zijian Chen, Wei Sun, Haoning Wu, Zicheng Zhang, Jun Jia, Zhongpeng Ji, Fengyu Sun, Shangling Jui, Xiongkuo Min, Guangtao Zhai, and Wenjun Zhang. 2023. Exploring the Naturalness of AI-Generated Images. https://doi.org/10.48550/ARXIV.2312.05476 Version Number: 3.
- Cutzu et al. (2003) F. Cutzu, R. Hammoud, and A. Leykin. 2003. Estimating the photorealism of images: distinguishing paintings from photographs. In 2003 IEEE Computer Society Conference on Computer Vision and Pattern Recognition, 2003. Proceedings., Vol. 2. IEEE Comput. Soc, Madison, WI, USA, II–305–12. https://doi.org/10.1109/CVPR.2003.1211484
- De Ridder (1996) Huib De Ridder. 1996. Naturalness and Image Quality: Saturation and Lightness Variation in Color Images of Natural Scenes. Journal of Imaging Science and Technology 40, 6 (Nov. 1996), 487–493. https://doi.org/10.2352/J.ImagingSci.Technol.1996.40.6.art00003
- Designboom (2023) Designboom. 2023. Puffy pope, Trump arrest, blue plague: How fake AI-generated images fooled the internet. Designboom. https://www.designboom.com/technology/puffy-pope-trump-arrest-blue-plague-fake-ai-03-29-2023/ Accessed: 2024-09-07.
- Devlin and Cheetham (2023) Kayleen Devlin and Joshua Cheetham. 2023. Fake Trump arrest photos: How to spot an AI-generated image. BBC. https://www.bbc.com/news/world-us-canada-65069316
- DiResta and Goldstein (2024) Renee DiResta and Josh A Goldstein. 2024. How spammers and scammers leverage ai-generated images on facebook for audience growth. arXiv preprint arXiv:2403.12838 (2024).
- Fan et al. (2020) Shaojing Fan, Bryan L. Koenig, Qi Zhao, and Mohan S. Kankanhalli. 2020. A Deeper Look at Human Visual Perception of Images. SN Computer Science 1, 1 (Jan. 2020), 58. https://doi.org/10.1007/s42979-019-0061-5
- Fan et al. (2014) Shaojing Fan, Tian-Tsong Ng, Jonathan S. Herberg, Bryan L. Koenig, Cheston Y.-C. Tan, and Rangding Wang. 2014. An Automated Estimator of Image Visual Realism Based on Human Cognition. In 2014 IEEE Conference on Computer Vision and Pattern Recognition. IEEE, Columbus, OH, USA, 4201–4208. https://doi.org/10.1109/CVPR.2014.535
- Ferwerda (2003) James A. Ferwerda. 2003. Three varieties of realism in computer graphics. In Proceedings of SPIE: Human Vision and Electronic Imaging VIII, Bernice E. Rogowitz and Thrasyvoulos N. Pappas (Eds.). SPIE, Santa Clara, CA, 290. https://doi.org/10.1117/12.473899
- Germain (1974) Edward Germain. 1974. Four Surrealist Images. Visible Language 8, 4 (1974), 319–332. https://journals.uc.edu/index.php/vl/article/view/5173
- Gupta et al. (2013) Aditi Gupta, Hemank Lamba, Ponnurangam Kumaraguru, and Anupam Joshi. 2013. Faking Sandy: Characterizing and Identifying Fake Images on Twitter during Hurricane Sandy. In Proceedings of the 22nd International Conference on World Wide Web (WWW 2013 Companion). ACM, International World Wide Web Conferences Steering Committee, Rio de Janeiro, Brazil, 729–736.
- Hausken (2024) Liv Hausken. 2024. Photorealism versus photography. AI-generated depiction in the age of visual disinformation. Journal of Aesthetics & Culture 16, 1 (Dec. 2024), 2340787. https://doi.org/10.1080/20004214.2024.2340787
- Huang (1997) Zhexue Huang. 1997. CLUSTERING LARGE DATA SETS WITH MIXED NUMERIC AND CATEGORICAL VALUES. In Proceedings of the 1st International Conference on Knowledge Discovery and Data Mining (KDD’97). AAAI Press, San Diego, CA, USA, 21–34. https://api.semanticscholar.org/CorpusID:3007488
- Iigaya et al. (2021) Kiyohito Iigaya, Sanghyun Yi, Iman A. Wahle, Koranis Tanwisuth, and John P. O’Doherty. 2021. Aesthetic preference for art can be predicted from a mixture of low- and high-level visual features. Nature Human Behaviour 5, 6 (May 2021), 743–755. https://doi.org/10.1038/s41562-021-01124-6
- Instagram (2024) Instagram. 2024. https://help.instagram.com/761121959519495 Accessed: 2024-09-07.
- Itti and Koch (2000) Laurent Itti and Christof Koch. 2000. A saliency-based search mechanism for overt and covert shifts of visual attention. Vision research 40, 10-12 (2000), 1489–1506.
- Joo et al. (2014) Jungseock Joo, Weixin Li, Francis F. Steen, and Song-Chun Zhu. 2014. Visual Persuasion: Inferring Communicative Intents of Images. In 2014 IEEE Conference on Computer Vision and Pattern Recognition. IEEE, Columbus, OH, USA, 216–223. https://doi.org/10.1109/CVPR.2014.35
- Joo and Steinert-Threlkeld (2022) Jungseock Joo and Zachary C. Steinert-Threlkeld. 2022. Image as Data: Automated Content Analysis for Visual Presentations of Political Actors and Events. Computational Communication Research 4, 1 (Feb. 2022), 1–26. https://doi.org/10.5117/CCR2022.1.001.JOO
- Kanuri et al. (2024) Vamsi K. Kanuri, Christian Hughes, and Brady T. Hodges. 2024. Standing out from the crowd: When and why color complexity in social media images increases user engagement. International Journal of Research in Marketing 41, 2 (June 2024), 174–193. https://doi.org/10.1016/j.ijresmar.2023.08.007
- Kasra et al. (2018) Mona Kasra, Cuihua Shen, and James F. O’Brien. 2018. Seeing Is Believing: How People Fail to Identify Fake Images on the Web. In Extended Abstracts of the 2018 CHI Conference on Human Factors in Computing Systems. ACM, Montreal QC Canada, 1–6. https://doi.org/10.1145/3170427.3188604
- Lee et al. (2013) Cha Lee, Gustavo A. Rincon, Greg Meyer, Tobias Hollerer, and Doug A. Bowman. 2013. The Effects of Visual Realism on Search Tasks in Mixed Reality Simulation. IEEE Transactions on Visualization and Computer Graphics 19, 4 (April 2013), 547–556. https://doi.org/10.1109/TVCG.2013.41
- Lehmuskallio et al. (2019) Asko Lehmuskallio, Jukka Häkkinen, and Janne Seppänen. 2019. Photorealistic computer-generated images are difficult to distinguish from digital photographs: a case study with professional photographers and photo-editors. Visual Communication 18, 4 (Nov. 2019), 427–451. https://doi.org/10.1177/1470357218759809
- Levine (2014) Timothy R Levine. 2014. Truth-default theory (TDT) a theory of human deception and deception detection. Journal of Language and Social Psychology 33, 4 (2014), 378–392.
- Li and Xie (2020) Yiyi Li and Ying Xie. 2020. Is a Picture Worth a Thousand Words? An Empirical Study of Image Content and Social Media Engagement. Journal of Marketing Research 57, 1 (Feb. 2020), 1–19. https://doi.org/10.1177/0022243719881113
- Lu and (Cindy) Yingdan Lu and Cuihua (Cindy) Shen. 2023. Unpacking Multimodal Fact-Checking: Features and Engagement of Fact-Checking Videos on Chinese TikTok (Douyin). Social Media + Society 9, 1 (Jan. 2023), 205630512211504. https://doi.org/10.1177/20563051221150406
- Lu and Peng (2024) Yingdan Lu and Yilang Peng. 2024. The Mobilizing Power of Visual Media Across Stages of Social-Mediated Protests. Political Communication 41, 4 (July 2024), 531–558. https://doi.org/10.1080/10584609.2024.2317951
- Lu et al. (2023) Zeyu Lu, Di Huang, LEI BAI, Jingjing Qu, Chengyue Wu, Xihui Liu, and Wanli Ouyang. 2023. Seeing is not always believing: Benchmarking Human and Model Perception of AI-Generated Images. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., New Orleans, LA, USA, 25435–25447. https://proceedings.neurips.cc/paper_files/paper/2023/file/505df5ea30f630661074145149274af0-Paper-Datasets_and_Benchmarks.pdf
- Mackenzie (2024) Jean Mackenzie. 2024. South Korea faces Deepfake Porn “emergency”. https://www.bbc.com/news/articles/cg4yerrg451o
- Malcolm (2024) Kim Malcolm. 2024. AI images and conspiracy theories are driving a push for Media Literacy Education. https://www.npr.org/2024/03/21/1239693671/ai-images-and-conspiracy-theories-are-driving-a-push-for-media-literacy-educatio
- Matz et al. (2019) Sandra C. Matz, Cristina Segalin, David Stillwell, Sandrine R. Müller, and Maarten W. Bos. 2019. Predicting the Personal Appeal of Marketing Images Using Computational Methods. Journal of Consumer Psychology 29, 3 (July 2019), 370–390. https://doi.org/10.1002/jcpy.1092
- Muñoz and Towner (2017) Caroline Lego Muñoz and Terri L. Towner. 2017. The Image is the Message: Instagram Marketing and the 2016 Presidential Primary Season. Journal of Political Marketing 16, 3-4 (Oct. 2017), 290–318. https://doi.org/10.1080/15377857.2017.1334254
- Neumüller (2018) Moritz Neumüller (Ed.). 2018. The Routledge Companion to Photography and Visual Culture (1 ed.). Routledge, London; New York. https://doi.org/10.4324/9781138604391
- Peng (2021) Yilang Peng. 2021. What Makes Politicians’ Instagram Posts Popular? Analyzing Social Media Strategies of Candidates and Office Holders with Computer Vision. The International Journal of Press/Politics 26, 1 (Jan. 2021), 143–166. https://doi.org/10.1177/1940161220964769
- Peng (2022) Yilang Peng. 2022. Athec: A Python Library for Computational Aesthetic Analysis of Visual Media in Social Science Research. Computational Communication Research 4, 1 (Feb. 2022), 35–50. https://doi.org/10.5117/CCR2022.1.009.PENG
- Peng et al. (2023a) Yilang Peng, Yingdan Lu, and Cuihua Shen. 2023a. An Agenda for Studying Credibility Perceptions of Visual Misinformation. Political Communication 40, 2 (March 2023), 225–237. https://doi.org/10.1080/10584609.2023.2175398
- Peng et al. (2023b) Yilang Peng, Tian Yang, and Kecheng Fang. 2023b. The dark side of entertainment? How viral entertaining media build an attention base for the far-right politics of The Epoch Times. new media & society 0, 0 (2023), 14614448231205893.
- Pereira et al. (2021) Lidiane Pereira, Wellingston C. Roberti Junior, and Rodrigo L. S. Silva. 2021. Photorealism in Mixed Reality: A Systematic Literature Review. International Journal of Virtual Reality 21, 1 (May 2021), 15–29. https://doi.org/10.20870/IJVR.2021.21.1.3166
- PetaPixel (2023) PetaPixel. 2023. AI Image Fools Judges and Wins Photography Contest. https://petapixel.com/2023/02/10/ai-image-fools-judges-and-wins-photography-contest/ Accessed: 2024-09-07.
- Pew Research Center (2023) Pew Research Center. 2023. Teens, Social Media and Technology 2023. https://www.pewresearch.org/internet/2023/12/11/teens-social-media-and-technology-2023/ Accessed: 2024-09-07.
- Pieters (2010) Cornelis Pieter Pieters. 2010. Into complexity: a pattern-oriented approach to stakeholder communications = orgaan in complexiteit : een patroon-oriëntatie. s.n., S.l. OCLC: 700747632.
- Qian et al. (2024) Sijia Qian, Yingdan Lu, Yilang Peng, Cuihua (Cindy) Shen, and Huacen Xu. 2024. Convergence or divergence? A cross-platform analysis of climate change visual content categories, features, and social media engagement on Twitter and Instagram. Public Relations Review 50, 2 (June 2024), 102454. https://doi.org/10.1016/j.pubrev.2024.102454
- Rademacher et al. (2001) Paul Rademacher, Jed Lengyel, Edward Cutrell, and Turner Whitted. 2001. Measuring the Perception of Visual Realism in Images. In Rendering Techniques 2001, Steven J. Gortler and Karol Myszkowski (Eds.). Springer Vienna, Vienna, 235–247. https://doi.org/10.1007/978-3-7091-6242-2_22 Series Title: Eurographics.
- Royal United Services Institute (2023) Royal United Services Institute. 2023. It’s Time to Stop Debunking AI-Generated Lies and Start Identifying Truth. https://www.rusi.org/explore-our-research/publications/commentary/its-time-stop-debunking-ai-generated-lies-and-start-identifying-truth Accessed: 2024-09-07.
- Sharma and Peng (2024) Muna Sharma and Yilang Peng. 2024. How Visual Aesthetics and Calorie Density Predict Food Image Popularity on Instagram: A Computer Vision Analysis. Health Communication 39, 3 (Feb. 2024), 577–591. https://doi.org/10.1080/10410236.2023.2175635
- Shen et al. (2021) Cuihua Shen, Mona Kasra, and James O’Brien. 2021. This photograph has been altered: Testing the effectiveness of image forensic labeling on news image credibility. arXiv 1, 1 (2021), 1–15. https://doi.org/10.48550/ARXIV.2101.07951 Publisher: arXiv Version Number: 1.
- Shen et al. (2019) Cuihua Shen, Mona Kasra, Wenjing Pan, Grace A Bassett, Yining Malloch, and James F O’Brien. 2019. Fake images: The effects of source, intermediary, and digital media literacy on contextual assessment of image credibility online. New Media & Society 21, 2 (Feb. 2019), 438–463. https://doi.org/10.1177/1461444818799526
- Sundar and Shyam (2008) S. Shyam Sundar and Shyam. 2008. The MAIN Model: A Heuristic Approach to Understanding Technology Effects on Credibility. In Digital Media, Youth, and Credibility. The MIT Press, Cambridge, MA, 73–100.
- Teachers College, Columbia University (2023) Teachers College, Columbia University. 2023. Artificial Intelligence Optimists Concerned About Risks of Visual Misinformation. https://www.tc.columbia.edu/articles/2023/october/artificial-intelligence-optimists-concerned-about-risks-of-visual-misinformation/ Accessed: 2024-09-07.
- Wagg (1998) Stephen Wagg (Ed.). 1998. Because I tell a joke or two: comedy, politics, and social difference. Routledge, London New York.
- Wang et al. (2024) Jiarui Wang, Huiyu Duan, Jing Liu, Shi Chen, Xiongkuo Min, and Guangtao Zhai. 2024. AIGCIQA2023: A Large-Scale Image Quality Assessment Database for AI Generated Images: From the Perspectives of Quality, Authenticity and Correspondence. In Artificial Intelligence, Lu Fang, Jian Pei, Guangtao Zhai, and Ruiping Wang (Eds.). Vol. 14474. Springer Nature Singapore, Singapore, 46–57. https://doi.org/10.1007/978-981-99-9119-8_5 Series Title: Lecture Notes in Computer Science.
- Wang and Doube (2011) Norman Wang and Wendy Doube. 2011. How real is reality? A perceptually motivated system for quantifying visual realism in digital images. In 2011 International Conference on Multimedia and Signal Processing. IEEE, Guilin, China, 141–149. https://doi.org/10.1109/CMSP.2011.172
- Wang et al. (2021) Xin Wang, Jingna Jin, Wenbo Liu, Zhipeng Liu, and Tao Yin. 2021. Emotional processing of sadness and disgust evoked by disaster scenes. Brain and Behavior 11, 12 (2021), e02421. https://doi.org/10.1002/brb3.2421
Appendix A Appendix
A.1. Appendix A
| Variables | Twitter (%) | IG (%) | |
| Content features | N = 1,858 | N = 1,785 | |
| Human | 67.49 | 68.68 | 0.54 |
| Human-like | 10.06 | 7.45 | 7.44** |
| Non-human living beings | 28.15 | 26.55 | 1.08 |
| Non-living objects | 98.44 | 98.60 | 0.07 |
| Add on captions | 5.44 | 7.39 | 5.51* |
| Fictional captions | 2.05 | 2.13 | 0.00 |
| Other captions | 1.72 | 2.63 | 3.14 |
| Disaster | 2.37 | 1.34 | 4.66* |
| Surrealism combination | 58.23 | 60.78 | 2.35 |
| Human features | N = 1,254 | N = 1,226 | |
| Face | 73.37 | 70.96 | 1.66 |
| - Identity | 14.13 | 10.31 | 5.60* |
| - Politician | 22.31 | 36.67 | 4.73* |
| - Celebrity | 61.54 | 38.89 | 10.05** |
| Aesthetic features | N = 1,858 | N = 1,785 | |
| Colored images | 97.95 | 93.00 | 51.20*** |
| High image quality | 92.63 | 98.38 | 67.87*** |
| AI production features | N = 1,858 | N = 1,785 | |
| Production flaw | 8.13 | 8.68 | 0.30 |
| Watermark | 0.75 | 1.51 | 4.06* |
| Note: , , . | |||